# Robot Synesthesia: In-Hand Manipulation with Visuotactile Sensing

### v2

@@ Abstract
Executing contact-rich manipulation tasks necessitates the fusion of tactile and visual feedback. However, the distinct nature of these modalities poses significant challenges. In this paper, we introduce a system that leverages visual and tactile sensory inputs to enable dexterous in-hand manipulation. Specifically, we propose Robot Synesthesia, a novel point cloud-based tactile representation inspired by human tactile-visual synesthesia. This approach allows for the simultaneous and seamless integration of both sensory inputs, offering richer spatial information and facilitating better reasoning about robot actions. The method, trained in a simulated environment and then deployed to a real robot, is applicable to various in-hand object rotation tasks. Comprehensive ablations are performed on how the integration of vision and touch can improve reinforcement learning and Sim2Real performance. Our project page is available at https://yingyuan0414.github.io/visuotactile/.
@@
$1$ University of California San Diego, CA, USA
$2$ Tsinghua University, Beijing, China
$3$ University of Illinois Urbana-Champaign, IL, USA
$4$ University of California Berkeley, CA, USA
$5$ Dongguk University, Seoul, South Korea
@@ II Related Work
Dexterous Manipulation presents a wealth of opportunities for broad applications [4, 5, 6, 7, 8, 9]. It enables the execution of intricate manipulative tasks, such as sliding [10, 11], rolling [12, 13, 14, 15, 16], pivoting [17, 18], and regrasping [19, 20, 21]. Earlier methods addressing dexterous manipulation were grounded in classical control [22, 23]. However, these methods rely on expert-engineered dynamics models, which restricts their utility for more complex tasks. To overcome these limitations, recent research has leveraged deep model-free RL for dexterous manipulation [24, 25, 26, 27, 28]. Further enriching these advancements, imitation learning and combining common RL with demonstrations has led to higher sample efficiency and more natural manipulation behaviors [29, 30, 31, 32, 33, 34, 35, 36, 37]. Dexterous in-hand manipulation has been a focal point of research interest recently [38, 39, 40, 41, 42, 43]. To generalize to new objects, researchers have explored different sensors to capture object geometry and dynamic properties. Qi et al.[41] demonstrated that policy could infer object position and physical properties using proprioceptive history. But without explicit object information, it was only effective for z-axis rotation tasks. Yin et al. [42] proposed integrating binary tactile signals with proprioception for this task. However, the tactile signal was too sparse to capture detailed geometric attributes and thus could not handle objects with non-convex shapes. To solve this, Chen et al. [44] utilized depth information to aid object rotation while Guzey et al. [45] learned non-parametric policies on both vision and touch signal. Most similar to us, a recent study [46] utilizes RGB images from optical tactile sensors and depth images for in-hand rotation. However, it necessitated continuous contact between the object and tactile sensors, constraining it to smaller objects that can be rotated on the fingertips. In contrast, our work does not impose any specific requirements on the object’s initial location and can handle objects of diverse shapes and sizes. Furthermore, our tactile point cloud representation provides explicit 3D information about the object and tactile sensors’ location, while their models are based on 2D images. As a result, our policy can solve tasks requiring more complex 3D spatial reasoning, such as rotating two balls simultaneously.
@@ III-A System Setup
As is shown in Figure 1, our hardware setup consists of an XArm6 robot arm and a 16-DOF Allegro Hand with a depth camera. We attach 16 Force-Sensing Resistors (FSR) as tactile sensors to the palm and finger links of the robot hand as suggested by [42]. We gather the contact signal from each sensor, then binarize the measurement according to a predetermined threshold $\theta_{th}$ to abridge the Sim-to-Real gap. We use Isaac Gym[62] as the rigid body physics simulator. The simulation setup is visualized in Figure Robot Synesthesia: In-Hand Manipulation with Visuotactile Sensing. We simulate 16 contact sensors and calculate binary contact signals in the same way as in real.
@@ V Experiments
We answer these questions through an extensive case study on the benchmark problems.
@@ V-A2 Evaluation Metric
To evaluate the performance of a trained policy, we use the following metrics as suggested by [41].
@@ V-C Stage II: Imitation learning with different sensing capabilities
In this section, we distill teacher policies to visuotactile-based policies and perform ablation study for different sensing capabilities. As shown in Table II, Touch refers to binary contact, Cam refers to camera-based point clouds, Aug refers to augmented point clouds, and Syn refers to proposed tactile point clouds. For rotating objects of regular shapes, visual policies achieve similar dexterity to each other. However, when it comes to more challenging objects like multi-way wrenches and two balls, our visual-tactile synesthesia method outperforms all the baselines.
@@ V-D Real-world Deployment
We transfer the visuotactile policies to the real robot without any fine-tuning and test whether visual policies continue to provide benefits. The results are shown in Table III. We further evaluate our double-ball policy on different sets of real-world balls, shown in Table IV and Figure 7. Although visual policy might show comparable performance for simple geometry in simulation, the advantage of integrating vision and touch becomes more significant when deployed to the real world. These results highlight the low domain gap of our proposed tactile point cloud representation. A rudimentary concatenation of tactile signals and the extracted features of point clouds could increase the challenge for the policy to comprehend their underlying relationship. In contrast, our proposed visual-tactile synesthesia approach generally offers benefits. We also observe that visual policies tend to operate more cautiously, making occasional adjustments to nudge the object back to the palm center, while policies lacking visual perception tend to execute an almost fixed sequence of motion, irrespective of the object’s deviation or instances of it becoming stuck.
@@ V-E Qualitative Analysis: Visualization of PointNet intermediates
TABLE IV: Evaluation of our double-ball policy (CRA/TTF) on different sets of real-world balls. Each policy is tested for 5 episodes. Each trial lasts 60 seconds.
Obs Type
(CRA/TTF) Two Tomatoes Two Potatoes Two Golfballs Golfball-Tomato Golfball-Potato Tomato-Potato
Touch 11.6/20.8 9.2/23.0 29.8/41.6 35.4/50.2 6.8/23.2 7.4/17.0
Cam+Aug 9.2/22.4 9.2/22.8 16.6/25.8 13.8/29.6 8.0/17.0 5.6/16.2
Touch+Cam+Aug 18.6/31.2 6.6/15.0 35.0/51.0 32.0/53.0 22.8/41.4 9.4/26.8
Touch+Cam+Aug+Syn 24.2/40.0 22.4/42.0 41.6/60.0 42.0/60.0 8.2/16.2 9.8/21.0
Figure 7: Visualization of our double-ball policy in the real-world testing on Tomato-Potato, Golfball-Tomato, and Two Golfballs.

### v3

@@ Abstract
Executing contact-rich manipulation tasks necessitates the fusion of tactile and visual feedback. However, the distinct nature of these modalities poses significant challenges. In this paper, we introduce a system that leverages visual and tactile sensory inputs to enable dexterous in-hand manipulation. Specifically, we propose Robot Synesthesia, a novel point cloud-based tactile representation inspired by human tactile-visual synesthesia. This approach allows for the simultaneous and seamless integration of both sensory inputs, offering richer spatial information and facilitating better reasoning about robot actions. Comprehensive ablations are performed on how the integration of vision and touch can improve reinforcement learning and Sim2Real performance. Our project page is available at https://yingyuan0414.github.io/visuotactile/.
@@
$1$ University of California San Diego
$2$ Tsinghua University
$3$ University of Illinois Urbana-Champaign
$4$ University of California Berkeley
$5$ Dongguk University
@@ II Related Work
Dexterous Manipulation presents a wealth of opportunities for broad applications [4, 5, 6, 7, 8, 9]. It enables the execution of intricate manipulative tasks, such as sliding [10, 11], rolling [12, 13, 14, 15, 16], pivoting [17, 18], and regrasping [19, 20, 21]. Earlier methods addressing dexterous manipulation were grounded in classical control [22, 23]. However, these methods rely on expert-engineered dynamics models, which restricts their utility for more complex tasks. To overcome these limitations, recent research has leveraged deep model-free RL for dexterous manipulation [24, 25, 26, 27, 28]. Further enriching these advancements, imitation learning and combining common RL with demonstrations has led to higher sample efficiency and more natural manipulation behaviors [29, 30, 31, 32, 33, 34, 35, 36, 37]. Dexterous in-hand manipulation has been a focal point of research interest recently [38, 39, 40, 41, 42, extreme, 43]. To generalize to new objects, researchers have explored different sensors to capture object geometry and dynamic properties. Qi et al.[41] demonstrated that policy could infer object position and physical properties using proprioceptive history. But without explicit object information, it was only effective for z-axis rotation tasks. Yin et al. [42] proposed integrating binary tactile signals with proprioception for this task. However, the tactile signal was too sparse to capture detailed geometric attributes and thus could not handle objects with non-convex shapes. To solve this, Chen et al. [44] utilized depth information to aid object rotation while Guzey et al. [45] learned non-parametric policies on both vision and touch signal. Most similar to us, a recent study [46] utilizes RGB images from optical tactile sensors and depth images for in-hand rotation. However, it necessitated continuous contact between the object and tactile sensors, constraining it to smaller objects that can be rotated on the fingertips. In contrast, our work does not impose any specific requirements on the object’s initial location and can handle objects of diverse shapes and sizes. Furthermore, our tactile point cloud representation provides explicit 3D information about the object and tactile sensors’ location, while their models are based on 2D images. As a result, our policy can solve tasks requiring more complex 3D spatial reasoning, such as rotating two balls simultaneously.
@@ III-A System Setup
As is shown in Figure 1, our hardware setup consists of an XArm6 robot arm and a 16-DOF Allegro Hand with a depth camera. We attach 16 Force-Sensing Resistors (FSR) as tactile sensors to the palm and finger links of the robot hand as suggested by [42]. We gather the contact signal from each sensor, then binarize the measurement according to a predetermined threshold $\theta_{th}$ to abridge the Sim-to-Real gap. We use Isaac Gym[62] as the rigid body physics simulator. The simulation setup is visualized in Figure Robot Synesthesia: In-Hand Manipulation with Visuotactile Sensing.
@@ V-A2 Evaluation Metric
To evaluate the policy performance, we use the following metrics as suggested by [41].
@@ V-C Stage II: Imitation learning with different sensing capabilities
In this stage, we distill teacher policies to visuotactile policies and perform ablation study for different sensing capabilities. As shown in Table II, Touch refers to binary contact, Cam refers to camera-based point clouds, Aug refers to augmented point clouds, and Syn refers to proposed tactile point clouds. For rotating objects of regular shapes, visual policies achieve similar dexterity to each other. However, when it comes to more challenging objects like multi-way wrenches and two balls, our method outperforms all the baselines.
@@ V-D Real-world Deployment
We transfer the visuotactile policies to the real robot without any fine-tuning and test whether visual policies continue to provide benefits. The results are shown in Table III. Although visual policy might show comparable performance for simple geometry in simulation, the advantage of integrating vision and touch becomes more significant when deployed to the real world. These results highlight the low domain gap of our proposed tactile point cloud representation. A rudimentary concatenation of tactile signals and the extracted features of point clouds could increase the challenge for the policy to comprehend their underlying relationship. In contrast, our proposed visual-tactile synesthesia approach generally offers benefits. We also observe that visual policies tend to operate more cautiously, making occasional adjustments to nudge the object back to the palm center, while policies lacking visual perception tend to execute an almost fixed sequence of motion, irrespective of the object’s deviation or instances of it becoming stuck.
@@ VI Conclusion
VII Acknowledgement
Acknowledgment: This work was supported, in part, by the Qualcomm Innovation Fellowship, and the Technology Innovation Program (20018112, Development of autonomous manipulation and gripping technology using imitation learning based on visual and tactile sensing) funded by the Ministry of Trade, Industry & Energy (MOTIE), Korea.